UT 22 Sep 2018 - Split D/E with SOUL Commissioning


Observers: B. Rothberg, O. Kuhn

OSA: S. Allanson

ISp: E. Solheaid

Mtn Manager: P. Hartley

UT 07:04 - currently closed for weather since UT 06:00, had an NFS mount failure, /newdata is inaccessible from any computer, and all IDL licenses are now locked out.

UT 07:48 - issues seem to be fixed, test images are in newdata and appear in modsDisp. Still closed for pea soup sky and high humidity.

UT 11:48 - i'm calling it, its yo-yo-ing too much with humidity - we are at -18, I don't have a quick program to do with MODS and I doubt I'd get a photon before we hit 95 again.

UT 11:49 - putting both MODS to sleep.

-- %USERSIG{BarryRothberg - 2018-09-22}%
Topic revision: r3 - 22 Sep 2018, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback